# Automated Genomic Interpretation via Concept Bottleneck Models for Medical Robotics

### v1

@@ IV-E Automated Decision Layer
We evaluate the final decision layer of our genetic interpretation module. This layer integrates predicted concepts, classifier confidence, and uncertainty into a calibrated scoring function, producing interpretable decisions that can be directly used in automated genetic analysis. We benchmark our decision head on two datasets (our gag set and LANL). For comparison, we use a rule-based proxy policy constructed from fixed thresholds on CpG, GC, and uncertainty. While the proxy provides weak supervision, it is rigid and often flags excessive retests. Our method learns from the same supervision but optimizes a cost-aware loss and applies temperature calibration. We report Accuracy, F1, AUROC, ECE (Expected Calibration Error), and Utility. RetestRate is also measured to evaluate cost–benefit trade-offs.
@@
Compared with the rule-based proxy, our decision head improves Utility (0.735 vs. 0.683 on our dataset; 0.724 vs. 0.676 on LANL) and reduces RetestRate (0.198 vs. 0.245; 0.212 vs. 0.268). This shows that the module achieves more favorable cost–benefit trade-offs, reducing unnecessary retests while preserving sensitivity. At the same time, AUROC remains above 0.91, and ECE below 0.05, confirming both strong discrimination and reliable calibration. The decision layer translates concept-level evidence into calibrated and cost-efficient decisions, reducing manual analysis burden and providing a ready-to-use component for clinical automation and medical robotics.
@@ IV-F Ablation Study
We ablate the automated genomic interpretation module by removing concept-fidelity supervision, prior-consistency, KL alignment, calibration, and by altering the concept pathway or the CGR front-end. We report concept fidelity ($R^{2}$, $r$), AUROC, calibration (ECE), and expected clinical cost.

### v2

@@ IV-E Automated Decision Layer
In Table V, we evaluate the final decision layer of our genetic interpretation module. This layer integrates predicted concepts, classifier confidence, and uncertainty into a calibrated scoring function, producing interpretable decisions that can be directly used in automated genetic analysis. We benchmark our decision head on two datasets (our gag set and LANL). For comparison, we use a rule-based proxy policy constructed from fixed thresholds on CpG, GC, and uncertainty. While the proxy provides weak supervision, it is rigid and often flags excessive retests. Our method learns from the same supervision but optimizes a cost-aware loss and applies temperature calibration. We report Accuracy, F1, AUROC, ECE (Expected Calibration Error), and Utility. RetestRate is also measured to evaluate cost–benefit trade-offs. Compared with the rule-based proxy, our decision head improves Utility (0.735 vs. 0.683 on our dataset; 0.724 vs. 0.676 on LANL) and reduces RetestRate (0.198 vs. 0.245; 0.212 vs. 0.268). This shows that the module achieves more favorable cost–benefit trade-offs, reducing unnecessary retests while preserving sensitivity. At the same time, AUROC remains above 0.91, and ECE below 0.05, confirming both strong discrimination and reliable calibration. The decision layer translates concept-level evidence into calibrated and cost-efficient decisions, reducing manual analysis burden and providing a ready-to-use component for clinical automation and medical robotics.
@@ IV-F Ablation Study
We ablate our module by removing concept-fidelity supervision, prior-consistency, KL alignment, calibration, and by altering the concept pathway or the CGR front-end. We report concept fidelity ($R^{2}$, $r$), AUROC, calibration (ECE), and expected clinical cost.